OpenStudy (anonymous):

The Celsius temperature equals five-ninths times the difference between the Fahrenheit temperature and 32. Find the Fahrenheit temperature when the Celsius temperature is 15°.

OpenStudy (lilg132):

have you written the question correctly?

OpenStudy (lilg132):

58.9999999784

OpenStudy (lilg132):

15 * 5/9 + 32 = farenheit 15 / 0.555555556 = 26.9999999784 + 32 = 58.9999999784 round it up 59 degrees fahrenheit dont forget to click good answer :)